OpenStudy (anonymous):

Two parallel chords in the same circle have lengths of 30cm and 48 cm. The circle has a radius of 25 cm. How far apart are the chords?

OpenStudy (anonymous):

would the answer be \[\approx 11?\]

OpenStudy (anonymous):

umm what were ur steps?

OpenStudy (anonymous):

I'm not sure....

OpenStudy (anonymous):

BA is the chord of length 30 cm and P is its midpoint. Triangle APC is a right triangle, the length of CA is 25 cm and the length of AP is 30/2 = 15 cm.

OpenStudy (anonymous):

hadn't thought thay they could be on the same side of the radius.. And the textbook doesn't specify. I know the answers are in the back, but they don't have processes to find them.

OpenStudy (dumbcow):

They are 13 cm apart

Directrix (directrix):

 Distance is perpendicular distance. The perpendicular to a chord bisects the chord. In each case, we get right triangles. Pythagorean Triples come in handy as (15, 20, 25) is a 5 multiple of the (3,4,5) triple and (7, 24,25) is a primitive triple. Regardless, the Pythagorean Theorem can be used: 25^2 = 15^2 + h^2 where h = 20 and then again as 25^2 = 24^2 + d^2 and d = 7. 7 + 20 = 27 for the distance between the parallel chords.